OpenStudy (anonymous):

Math Analysis Find the indicated term of the geometric sequence t1=-5, r=1/4. Find t4 :)

OpenStudy (campbell_st):

are you sure its and arithmetic series... as you have given a letter r which stands for the common ratio in a geometric series

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

sorry it's geometric sequence xD

OpenStudy (campbell_st):

then a term in a geometric series is \[T _{n} = ar^{n -1}\] you have a = -5, r = 1/4 and n = 4 just substitute and evaluate \[T _{4} = -5 \times (\frac{1}{4})^{ 4-1}\]

OpenStudy (anonymous):

ok, ty <33